2x-y+4z=19
-x+3y-2z=-7
4x+2y+3y=37

Answer by stanbon(48536)

2x-y+4z=19
-x+3y-2z=-7
4x+2y+3y=37
---------
Solved using matrices:
X=8
Y=1
Z=1
Cheers,
Stan H.